Christopher is 4 times as old as Nadia and is also 30 years older than Nadia
Answer by ankor@dixie-net.com(22740)

Let c = C's age
let n = N's age
:'
Write an equation for each statement
:
Christopher is 4 times as old as Nadia
c = 4n
:
and is also 30 years older than Nadia
c = n + 30
Replace c with 4n
4n = n + 30
4n - n = 30
3n = 30
n = 30/3
n = 10 yrs old
then
4(10) = 40 yrs is C's age